OpenStudy (anonymous):

Solve the system of equations. y = 3x + 1 4y = 12x + 4

OpenStudy (anonymous):

write out your equations like this: 4y= 12x+4 y-3x=1. You see what I did? I subtracted the 3x so the variables could line up

OpenStudy (fellowroot):

4y=12x+4

OpenStudy (anonymous):

oops sorry i messed up

OpenStudy (fellowroot):

its the same equation

OpenStudy (anonymous):

y-3x=1 4y-12x=4

OpenStudy (anonymous):

now variables are lined up

OpenStudy (anonymous):

-1(y-3x=1). To cancel one of the variables out, make it a negative. In this case, I chose the y by itself, so multiply the whole equation by -1. Then you get, -y+3x=-1

OpenStudy (anonymous):

4y-12x=4 -y+3x=-1.. Can you answer it from here?

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

I guess that didn't do anything

OpenStudy (anonymous):

what do i do if all the variable are not the same

OpenStudy (anonymous):

wait i think i got it

OpenStudy (anonymous):

quick question. Are you looking for the x or y?

OpenStudy (anonymous):

it just says solve the system of equations

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

welll if you get y by itself, you end up getting the equation, y=3x+1

OpenStudy (anonymous):

I have to go now.. sorry! You can ask someone else for help!

OpenStudy (anonymous):

can you not help anymore??

OpenStudy (anonymous):

it is the same equation it has infinite many solution

OpenStudy (anonymous):

both equations are the same! the 2nd one is just the 1st eq multiplied by 4.

OpenStudy (anonymous):

y-3x=1 4y-12x=4 1(y-3x=1)

OpenStudy (anonymous):

this is what i have

OpenStudy (anonymous):

I'm back. Like I said, both end up being the same exact question... y=3x+1, which means they have infinite solutions. I kinda forgot systems.. But yeah I guess that's my answer

OpenStudy (anonymous):

And ignore what I did with all that math

OpenStudy (anonymous):

oh ok..so no solution

OpenStudy (anonymous):

@coolaidd there is solution just let any of the variable arbitrary let y=t then solveequation 1 for x x=(t-1)/3 si solution is x=t/3-1/3 y=t it means it can have infinite many solutions whatever you want can take value of t . i hope it is clear